# Identification of Single-Treatment Effects in Factorial Experiments

### v2

@@ Appendix D Partial identification results
This result is due to the solution of the linear program detailed in the linear program detailed below. Because that’s the solution for the program, those are also the sharp bounds for the problem (cannot be improved upon). ∎
@@
This result is due to the solution of the linear program detailed in the linear program detailed below. Because that’s the solution for the program, those are also the sharp bounds for the problem (cannot be improved upon). ∎
@@
This result is due to the solution of the linear program detailed in the linear program detailed below. Because that’s the solution for the program, those are also the sharp bounds for the problem (cannot be improved upon). ∎
@@
This result is due to the solution of the linear program detailed in the linear program detailed below. Because that’s the solution for the program, those are also the sharp bounds for the problem (cannot be improved upon). ∎
@@ Appendix E Linear programs
This is the linear program for deriving sharp bounds for both estimands $\Pr(Y_{a_{1}}=y_{1})$ and the ATE using data on $\Pr(Y,A,B)$ and $\Pr(Y_{a,b})$. Each probabilistic quantity is transformed into principal strata, which will become variables of the program (Duarte et al.,, 2023). Data and axioms of probability become constraints, and the estimand, the objective function. The program might be solved using a symbolic or numeric linear program solver.
  • Variables: $[a_{0}b_{0}y_{0000},...,a_{1}b_{1}y_{1111}]$

### v3

@@ Appendix D Partial identification results
This result follows from the solution of the linear program detailed below. Because that solution is exact, these are the sharp bounds for the problem. ∎
@@
This result follows from the solution of the linear program detailed below. Because that solution is exact, these are the sharp bounds for the problem. ∎
@@
This result follows from the solution of the linear program detailed below. Because that solution is exact, these are the sharp bounds for the problem. ∎
@@
This result follows from the solution of the linear program detailed below. Because that solution is exact, these are the sharp bounds for the problem. ∎
@@ Appendix E Linear programs
This is the linear program for deriving sharp bounds for both estimands $\Pr(Y_{a_{1}}=y_{1})$ and the ATE using data on $\Pr(Y,A,B)$ and $\Pr(Y_{a,b})$. Each probabilistic quantity is transformed into principal strata, which will become variables of the program (Duarte et al., 2023). Data and axioms of probability become constraints, and the estimand, the objective function. The program might be solved using a symbolic or numeric linear program solver.
  • Variables: $a_{0}b_{0}y_{0000},...,a_{1}b_{1}y_{1111}]$